OpenStudy (anonymous):

x + 5y = -2 2x + y = 5 The point of intersection of the lines has a y-coordinate of

OpenStudy (anonymous):

To find the point of intersection of the 2 lines you need to solve the simultaneous equations: \[x + 5y = -2\] \[ 2x + y = 5\] Try taking 2 lots of the first equation away from the second, and see what you get.

OpenStudy (anonymous):

Substitute y = 5 - 2x in x +5y = 2 -> x+25-10x= -2